Determine the ΔS system given the reaction : 

       C6H12O6 (s) + O2(g) --------------→ CO2 (g) + H2O(l)


1
Expert's answer
2022-06-09T11:21:05-0400

∆S (C6H12O6 ) = 209,2 J/mol•K

∆S (O2 ) = 205,15 J/mol•K

∆S (CO2 ) = 213.79 J/mol•K

∆S (H2O) = 69.9 J/mol•K


C6H12O6 + 6 O2 = 6 CO2 + 6 H2O

∆S = Sproduced – Sreacted

∆S=(213,79*6 +69,9*6) - (205,15*6 - 209,2)=

=262,04 J/mol•K


Answer: ∆S= 262,04 J/mol•K